# Breaking the Bulkhead: Demystifying Cross-Namespace Reference Vulnerabilities in Kubernetes Operators

### v1

@@ 8. Conclusion
In this paper, we presented the first research on the security issues of Kubernetes Operators. We introduced Cross-Namespace Reference Vulnerability with two strategies, demonstrating how an attacker can bypass namespace isolation. We designed and implemented a static analysis suite to detect such vulnerabilities. Large-scale measurements illustrated that over 14% of Operators in the wild are potentially vulnerable. Our findings have been reported to the relevant developers, resulting in 7 confirmations and 6 CVEs by the time of submission, highlighting the critical need for enhanced security practices in Kubernetes Operators. We proposed possible mitigation solutions and open-sourced the static analysis suite to benefit the ecosystem.

### v2

@@ VIII Conclusion
In this paper, we presented the first in-depth research on Kubernetes Operator attacks, unveiling a long-neglected Cross-Namespace Reference Vulnerability with two strategies, demonstrating how an attacker can bypass namespace isolation. We designed and implemented a static analysis suite to conduct large-scale measurements, illustrating that over 14% of Operators in the wild are potentially vulnerable. Our findings have been reported to the relevant developers, resulting in 8 confirmations and 7 CVEs by the time of submission, highlighting the critical need for enhanced security practices in Kubernetes Operators. We proposed concrete mitigation solutions and open-sourced our code to benefit the ecosystem.
Ethics Considerations
We conducted our research with strict adherence to ethical guidelines. To collect Operator projects, we followed GitHub’s rate limits and usage policies. We validated vulnerabilities on our own Kubernetes clusters, ensuring that no third-party users or environments were affected. We responsibly disclosed vulnerabilities to all affected vendors. Thus, each affected project has at least 90 days to fix before publication. All vulnerability disclosures complied with the security policies of the respective vendors. All vulnerabilities in Table VI were validated and then reported to the vendors according to their vulnerability policies. To facilitate mitigation, we suggested available approaches, released mitigation samples and a lightweight mitigation generator, enabling quick adoption for practitioners. We will ensure all vulnerabilities discussed in the case studies are fixed or authorized to mention by the time of the final publication. We are actively coordinating with vendors and will publish full datasets at the proper time (expected conference date) to ensure developers have enough time to fix and minimize impacts.
Acknowledgment
We thank our reviewers and shepherd for their valuable feedback and comments. For authors, Ziyi Guo is supported by Google PhD Fellowship. Zhenyuan Li is supported by the National Natural Science Foundation of China under 62402419 and by the CCF-Tencent Rhino-Bird Open Research Fund. Any opinions, findings, and conclusions or recommendations expressed in this work are those of the author(s) and do not necessarily reflect the views of the institutions above.